OpenStudy (anonymous):

What is the exact value of the expression 8 sqrt 15 + 5 sqrt 5 + sqrt 15

OpenStudy (mertsj):

\[8\sqrt{15}+5\sqrt{5}+1\sqrt{15}\]

OpenStudy (mertsj):

Add the coefficients of the radicals that have the same radicand.

OpenStudy (anonymous):

yes, this is it

OpenStudy (anonymous):

i still dont get it

OpenStudy (anonymous):

do i start by \[8\sqrt{5} . 3 + 5\sqrt{5 } + 1\sqrt{5} . 3\]

OpenStudy (lgbasallote):

think of sqrt 15 as x \[8x + \sqrt 5 + x\] you can combine something right?